### v1

@@ 3.2 LongBench
For LongBench dataset, the LagKV method performs very well across different ratios and lag sizes. When $r=2\times$, for different $L$, it shows nearly zero loss compared to the baseline in both models. When $L=1024,r=8\times$, the method still retains approximate $90\%$ of the baseline performance. Since the compression ratio will increase when $L$ decreases, the worse case is $L=128,r=8\times$ for both models but the method maintains at least $85\%$ of the baseline performance.
@@ 3.3 Passkey Retrieval
The 64-digit passkey retrieval task is a challenging one for most token eviction strategies. As discussed in Yuan et al. (2024), the most succesful eviction strategy $H_{2}O$ Zhang et al. (2024) performs well in 7-digit task (scoring $100\%$ for all compression ratios) but degrades a lot in the 64-digit one (scoring $35\%$ for 4×4$\times$4 × in Llama-3) since the first token leakage issue. However, our method performs very well when $rL$ is sufficient large enough (scoring $96.57\%$ when $L=1024,r=4\times$). Our recursive compression strategy will not perform well for the setups with small $rL$ due to the fact that digitals usually require more tokens to be represented than the same length words. As shown in Fig. 2, the Qwen model which uses one token for one digit degenerates faster than the Llama model which represents three digits by one token with higher compression ratios. It hints us that we must choose the compression ratio and the lag size carefully in considering the length of the expected content.  A.1 shows all the details of the needle results.
@@
Figure 2: The needle score vs different setups of rL. The horizontal dash-dot line is the baseline for each model. The x-axis is in log scale. We put two vertical lines: $x=64$ (solid blue) and $x=128$ (green dash) for guidelines.
@@ 4 Related Works
The $L_{2}$ Norm-Based KV compression  Devoto et al. (2024) is an existing eviction approach that relies solely on KV information to compress the KV cache. This method computes token scores using the negative norm of key states. However, unlike our approach, it does not employ recursive compression during the prefill stage. As discussed in A.2, this method shows limitations in the 64-digit passkey retrieval task when the method is adapted to a recursive framework.
@@ 5 Conclusion
Unlike existing approaches, LagKV employs a recursive, attention-weight-free strategy in both prefill and decode stages to determine token importance for future processing. It’s indepent from query states and the rest part of the long prompt. Therefore our method offers a novel perspective on LLM mechanisms, shedding light on their inner workings in a fundamentally different way.
@@ A.1 Detail Rresults of Passkey Retrieval
Here, we present all the Needle-in-a-Haystack results with 64-digit Passkey Retrieval for different setups in Fig.3 and 4.
@@
Figure 3: The 64-digit Passkey Retrieval of Llama-3.1-8B-Instruct for different setups.
@@
Figure 4: The 64-digit Passkey Retrieval of Qwen-2.5-7B-Instruct for different setups.
A.2 Variants From LagKV Framework
Here we present two variants from LagKV. Both of them will only change the scoring methods but keep the attention sink and sliding window unchanged. And we only use the 64-digits passkey retrieval task which can easily distinguish eviction strategies as the detector. Among these tests, we keep $S=16,L=1024$ as constant.
The first one is called LocalKV which only skips using the reference from the next joint chunk tokens but replacing the equation Eq. 5 and 6 by the following equations:
$min_{i}^{p,K,V}=min_{seq}({K_{i}^{p},V_{i}^{p}})$ (12)
$max_{i}^{p,K,V}=max_{seq}({K_{i}^{p},V_{i}^{p}})$ (13)
$score_{i}=-Norm(K_{i})$ (14)
The results of the 64-digits passkey retrieval task is present in Fig. 5. As we can see, the LagKV method is always the best one especially in the high compression ratios. The $L_{2}$ norm method shows very limited performance.
Llama-3.1-8B-Instruct
Qwen-2.5-7B-Instruct
Figure 5: The 64-digit Passkey Retrieval scores of different variants and compression ratios.

### v2

@@ 4.1 LongBench
For the LongBench dataset, the method performs very well across different ratios and lag sizes. When $L=1024,r=8\times$, the method still retains approximate $90\%$ of the baseline performance. Since the compression ratio will increase when $L$ decreases, the worse case is $L=128,r=8\times$ for both models but the method maintains at least $85\%$ of the baseline performance.
@@ 4.2 Passkey Retrieval
The 64-digit passkey retrieval task is a challenging one for most token eviction strategies. As discussed in Yuan et al. (2024), the most successful eviction strategy $H_{2}O$ Zhang et al. (2024) performs well in 7-digit task (scoring $100\%$ for all compression ratios) but degrades a lot in the 64-digit one (for 4×4$\times$4 × in Llama-3, exact match score is $35\%$ and partial match score is $70.8\%$). It happens because the strategy applies its compression after the prefill is done which means the FGT is not affected by the compression and the 7-digit passkey usually takes only 2 or 3 tokens. When the passkey size increases to 64, much more generated tokens are impacted by the compression. Many token-evict algorithms are struggling to maintain their performance in this case. In contrast, our method performs very well when the product of $r$ and $L$ is sufficient large enough (for $L=1024,r=4\times$ in Llama model, exact math score is $89\%$ and partial match score is $96.57\%$).
Our recursive compression strategy will not perform well for the setups with small $rL$ due to the fact that when the recursive window size is compressed to be close to or less than the length of the queried content, it’s highly possible that only a small portion of the wanted information will be kept. In the task of 64-digit passkey retrieval, because digits usually require more tokens to be represented than the same length words, the number of expected tokens is much larger than the similar tasks in LongBench sub tasks like Document QA, that leads to its results are more sensitive to small $rL$. As shown in Fig. 2, the Qwen model which uses one token for one digit degenerates faster than the Llama model which represents three digits by one token with smaller $rL$. It hints us that we must choose the compression ratio and the lag size carefully in considering the length of the expected content and the tokenizer of the LLM.  A.1 shows all the details of the needle results.
(a) Llama-3.1-8B
(b) Qwen-2.5-7B
Figure 2: The needle score vs different setups of rL. The horizontal dash-dot line is the baseline for each model. The x-axis is in log scale. We put two vertical lines $rL=64$ (solid blue) and $rL=128$ (dash green) for guidelines.
4.3 Chunk-by-Chunk Compression in Prefill Stage
To enable chunk-by-chunk compression during prefill, we have to split the retrieval tokens like our recursive compression for long context with prefilling the first $S+2L$ tokens and then $L$ each time until all input tokens are prefilled. In such a way, the hidden values after the first chunk will be different from default prefill ones since less tokens are seen in the forwarding. Then, the FGT may be different too. With the chunk-by-chunk prefill compression, we calculated the FGT accuracy which is defined as the ratio of FGT same as the default prefill ones and also the overall needle scores, shown in Fig. 3.
(a) Llama-3.1-8B
(b) Qwen-2.5-7B
Figure 3: The needle score and FGT accuracy for different prefill methods with $L=1024$ only. The horizontal dash-dot line is the baseline for both needle scores and FGT accuracy since they are overlapping.
The chunked prefill definitely diminishes the FGT accuracy as it drops from $100\%$ to around $80\%$ for $r=8\times$ in both models. But we do not see it has a strong dependence on sequence lengths or needle depths in Fig. 4. These confirm that our method is able to retain the major part of the baseline capabilities in the case with long sequence hidden values impacted by the compression. It ensures that LagKV will deliver a good performance in the long generation scenarios.
$L=1024,r=2\times$
$L=1024,r=4\times$
$L=1024,r=6\times$
$L=1024,r=8\times$
@@ 4.3 Chunk-by-Chunk Compression in Prefill Stage
Figure 4: First Generated Token Accuracy for different setups, sequence lengths and needle depths with chunked prefill. It tests three trials on each depth.
Meanwhile, we also notice that the FGT accuracy and overall needle scores suffer more degradation in Llama model with chunked prefill. It is mainly because different models exhibit various abilities of stable long generation Quan et al. (2024).
4.4 Scoring Methods
We present two different scoring variants from LagKV. Both of them will only change the scoring methods but keep the attention sink and sliding window unchanged. And we only use the 64-digits passkey retrieval task which can easily distinguish eviction strategies as the detector. Among these tests, we keep $S=16,L=1024$ as constant.
The first one is called LocalKV which only skips using the reference from the next joint chunk tokens but replacing the equation Eq. 7 and 8 by the following equations:
$min_{i}^{p,Z}=min_{seq}({Z_{i}^{p}})$ (14)
$max_{i}^{p,Z}=max_{seq}({Z_{i}^{p}})$ (15)
$score_{i}=-Norm(K_{i})$ (16)
The results of the 64-digit passkey retrieval task are present in Fig. 5 and  6 with partial match scores and exact match scores. As we can see, the LagKV method is always the best one especially in the high compression ratios and the exact match cases. The LocalKV variant performs closely to LagKV at low compression ratios but degrades significantly at higher ones. This behavior stems from the similarity between local and remote max-min statistical values, which aligns with the token-wise locality.
Since the setup of $L=1024$ will have a chunk that fully covers the passkey when the context is shorter than 2K or the passkey is at $100\%$ depth, the bottom line of the exact match score will be about $27\%$ if the selected tokens did not mess up the output. That means the $L_{2}$ norm variant shows very limited performance with a constant exact match score $27\%$ for all compression ratios and models.
Llama-3.1-8B-Instruct
Qwen-2.5-7B-Instruct
Figure 5: The 64-digit Passkey Retrieval partial match scores of different variants and compression ratios.
Llama-3.1-8B-Instruct
Qwen-2.5-7B-Instruct
Figure 6: The 64-digit Passkey Retrieval exact match scores of different variants and compression ratios.
@@ 5 Related Works
The $L_{2}$ Norm-Based KV compression  Devoto et al. (2024) is an existing eviction approach that relies solely on KV information to compress the KV cache. This method computes token scores using the negative norm of key states. In contrast to our derivation from the autoregressive process and the token-wise locality, their method is formed by comparing the attention loss.
FINCH Corallo and Papotti (2024) introduces a prompt-guided KV compression method for the prefill stage, employing a chunk-by-chunk approach with instruction tokens appended to each document chunk. This design ensures the computation of attention submatrices between instructions and document chunks, enabling subsequent KV cache filtering. In contrast, our proposed chunked prefilling method operates without instructions, making it compatible with multi-turn queries. In other words, our approach transforms a causal LLM into a compressor capable of condensing long documents into compressed KV sequences, which can later be decompressed under varying instructions without reconstruction.
@@ 6 Conclusion
Unlike existing approaches, LagKV employs a recursive attention-weight-free strategy in both prefill and decode stages to determine token importance for future processing. It is independent from query states and the rest part of the long prompt. Therefore our method offers a novel perspective on LLM mechanisms, shedding light on their inner workings in a fundamentally different way.
@@ A.1 Detail Rresults of Passkey Retrieval
Here, we present all the Needle-in-a-Haystack results with 64-digit Passkey Retrieval for different setups. The partial matching results are in Fig.7 and 8 while Fig.9 and 10 with exact matching. Overall accuracies are noted within parentheses on the top-right corner of each sub graph.
@@
Figure 7: The 64-digit Passkey Retrieval of Llama-3.1-8B-Instruct for different setups with partial matching.
@@
Figure 8: The 64-digit Passkey Retrieval of Qwen-2.5-7B-Instruct for different setups with partial matching.
$L=1024,r=2\times$
$L=1024,r=4\times$
$L=1024,r=6\times$
$L=1024,r=8\times$
$L=512,r=2\times$
$L=512,r=4\times$
$L=512,r=6\times$
$L=512,r=8\times$
$L=128,r=2\times$
$L=128,r=4\times$
$L=128,r=6\times$
$L=128,r=8\times$
Figure 9: The 64-digit Passkey Retrieval of Llama-3.1-8B-Instruct for different setups with exact matching.
$L=1024,r=2\times$
$L=1024,r=4\times$
$L=1024,r=6\times$
$L=1024,r=8\times$
$L=512,r=2\times$
$L=512,r=4\times$
$L=512,r=6\times$
$L=512,r=8\times$
$L=128,r=2\times$
$L=128,r=4\times$
$L=128,r=6\times$
$L=128,r=8\times$
Figure 10: The 64-digit Passkey Retrieval of Qwen-2.5-7B-Instruct for different setups with exact matching.